You are analyzing the U.S. equity market based upon the S&P Industrials Index and using the present value of free cash flow to equity technique. Your inputs are as follows:

Beginning FCFE: $90
k = 0.09
Growth Rate:
Year 1–3: 9%
4–6: 8%
7 and beyond 7%
  1. Assuming that the current value for the S&P Industrials Index is 5,500, would you underweight, overweight, or market weight the U.S. equity market? Do not round intermediate calculations. Round your answer to the nearest cent.

    You should -Select-underweightoverweightmarket weightItem 1 the U.S. equity market as the estimated value of the stock of $   is -Select-higher thanlower thanequal torItem 3 the S&P Industrials Index.

  2. Assume that there is a 2 percent increase in the rate of inflation — what would be the market’s value, and how would you weight the U.S. market? Assume that the required return would increase from 9% to 11%, decreasing the value. Also assume that the nominal cash flow growth rates would increase for all time periods by two percentage points. Do not round intermediate calculations. Round your answer to the nearest cent.

    You should -Select-underweightoverweightmarket weightItem 4 the U.S. equity market as the estimated value of the stock of $   is -Select-higher thanlower thanequal torItem 6 than the S&P Industrials Index.
Answer #1
F0 Beginning FCFE $90
g1 Growth Rate Year1-3 9%
F1=F0*(1+g1) FCFE in Year 1 $98.10
F2=F1*(1+g1) FCFE in Year 2 $106.93
F3=F2*(1+g1) FCFE in Year 3 $116.55
g2 Growth Rate Year4-6 8%
F4=F3*(1+g2) FCFE in Year 4 $125.88
F5=F4*(1+g2) FCFE in Year 5 $135.95
F6=F5*(1+g2) FCFE in Year 6 $146.82
g3 Growth Rate Year7 and beyond 7%
F7=F6*(1+g3) FCFE in Year 7 $157.10
Present Value (PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=Discount Rate=k=0.09
N=Year of Cash Flow
Horizon Value in year 6 under constant growth rate of 7%
H6=F7/(k-g3) Horizon value in year 6 =157.10/(0.09-0.07) $7,855.02
N Year 1 2 3 4 5 6
a Free Cash flow $98.10 $106.93 $116.55 $125.88 $135.95 $146.82
b Horizon value $7,855.02
c=a+b TotalCash Flow $98.10 $106.93 $116.55 $125.88 $135.95 $8,001.84 SUM
PV=c/(1+0.09)^N Present value of Free Cash Flow to Equity $90.00 $90.00 $90.00 $89.17 $88.36 $4,771.23 $5,218.77
Total Present Value of Free cash flow to equity $5,218.77
You should underweight the U.S. equity market the S&P Industrials Index.
as the estimated value of the stock of $   is lower thanthe S&P industrial index
b
F0 Beginning FCFE $90
g1 Growth Rate Year1-3 11%
F1=F0*(1+g1) FCFE in Year 1 $99.90
F2=F1*(1+g1) FCFE in Year 2 $110.89
F3=F2*(1+g1) FCFE in Year 3 $123.09
g2 Growth Rate Year4-6 10%
F4=F3*(1+g2) FCFE in Year 4 $135.40
F5=F4*(1+g2) FCFE in Year 5 $148.94
F6=F5*(1+g2) FCFE in Year 6 $163.83
g3 Growth Rate Year7 and beyond 9%
F7=F6*(1+g3) FCFE in Year 7 $178.57

Growth Rate Year7 and beyond FCFE in Year 7 9% F7-F6*(1+g3 178.57 Present Value (PV) of Cash Flow: (Cash Flow)/((1+i)AN) i-Di